What is the length of the diagonal of a rectangle that has a width of 3 inches and a perimeter of 14 inches?
Answer by Adrish(44)

Solutions: Given: width=3inches
Perimeter=14 inches
Diagonal= Sqrt(length^2 + Width^2)
First we will find the length
P=2L + 2W
14=2L +2x3
14-6=2L
8=2L
L=8/2
L=4 inches
Now we will find the diagonal of the rectangle
Diagonal= sqrt(4^2 + 3^2)
D=sqrt(16 + 9)
D=Sqrt(25)
D=5